OpenStudy (anonymous):

What is the value of the range of the function f(x) = x² + 2 for the domain value One Quarter

OpenStudy (haseeb96):

f(0.25)=(0.25)^2+2 or f(1\4)=(1\4)^2+2 f(0.25)=0.0625+2 or f(1\4)= 1\8 +2 f(0.25)= 2.0625 or f(1\4)=1+16\2 or f(1\4)=17\2 2.0625 or 17\2 is the range of the f(0.25) or f(1\4) respectively this is the correct anwer